OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x - 2y = -4 3x + y = 12 A. -2y - 4 B. 2y - 4 C. 2y + 4 D. -2y + 4

OpenStudy (anonymous):

@DarkBlueChocobo @kirbykirby @TheRaggedyDoctor

OpenStudy (anonymous):

@Gabylovesyou @leonardo0430 @paki

OpenStudy (kirbykirby):

Basically what the question is asking is to isolate the x variable in the first equation, and once it is isolated, that will be the expression to substitute into the 2nd equation. So isolating x, we must keep it on one side of the equality: -x - 2y = -4 You can add 2y to both sides to get rid of it on the left side: -x -2y + 2y = -4 + 2y and simplify: -x = -4 + 2y Now... can you figure out how you might red "rid" of the negative in front of the x? :)